### v1

# Towards Robust Spatio-Temporal Auto-Regressive Prediction: Adams-Bashforth Time Integration with Adaptive Multi-Step Rollout

Abstract
This study addresses the critical challenge of error accumulation in spatio-temporal auto-regressive predictions within scientific machine learning models by introducing innovative temporal integration schemes and adaptive multi-step rollout strategies. We present a comprehensive analysis of time integration methods, highlighting the adaptation of the two-step Adams-Bashforth scheme to enhance long-term prediction robustness in auto-regressive models. Additionally, we improve temporal prediction accuracy through a multi-step rollout strategy that incorporates multiple future time steps during training, supported by three newly proposed approaches that dynamically adjust the importance of each future step. By integrating the Adams-Bashforth scheme with adaptive multi-step strategies, our graph neural network-based auto-regressive model accurately predicts 350 future time steps, even under practical constraints such as limited training data and minimal model capacity—achieving an error of only 1.6% compared to the vanilla auto-regressive approach. Moreover, our framework demonstrates an 83% improvement in rollout performance over the standard noise injection method, a standard technique for enhancing long-term rollout performance. Its effectiveness is further validated in more challenging scenarios with truncated meshes, showcasing its adaptability and robustness in practical applications. This work introduces a versatile framework for robust long-term spatio-temporal auto-regressive predictions, effectively mitigating error accumulation across various model types and engineering discipline.
keywords:
Spatio-temporal auto-regressive prediction , Scientific machine learning , Graph neural networks , Two-step Adams-Bashforth method , Adaptive multi-step rollout
journal: Applied Soft Computing
[…]
[inst1]organization=Cho Chun Shik Graduate School of Mobility, Korea Advanced Institute of Science and Technology,city=Daejeon, postcode=34051, country=Republic of Korea
[…]
[inst2]organization=FLOW, Engineering Mechanics, KTH Royal Institute of Technology,city=Stockholm, postcode=SE-100 44, country=Sweden
[…]
[inst3]organization=Narnia Labs,city=Daejeon, postcode=34051, country=Republic of Korea
1 Introduction
In the era of scientific machine learning (SciML), auto-regressive (AR) models have emerged as powerful tools for spatio-temporal prediction across various engineering disciplines, particularly in physical domains like fluid dynamics [1, 2, 3, 4]. They predict future states by recursively using their own previous predictions as inputs for subsequent forecasts—specifically, each new prediction becomes part of the input sequence for the next prediction step, creating a chain of sequential forecasts based on historical data. This recursive approach facilitates real-time forecasting and dynamic decision-making across a wide range of engineering applications [5, 6, 7]. Unlike other emerging SciML approaches such as physics-informed neural networks or DeepONet, which incorporate the time coordinate directly into the input and therefore violate temporal causality [8, 9, 10], AR models predict future states based on past historical information. This sequential prediction aligns with the natural progression of physical processes, preserving causality by ensuring that each snapshot at a given time depends solely on preceding snapshots, without any influence from future information.
However, AR models have inherent limitations, most notably error accumulation during long-term rollouts [2, 11, 12]. Since each prediction depends on the previous output, any inaccuracies introduced at one step can propagate and amplify in subsequent steps, leading to significant deviations from the true physical behavior over time. Recent approaches attempt to address this limitation by combining data-driven models with traditional numerical solvers, using the latter to recalibrate data-driven predictions when SciML model errors exceed certain thresholds [13, 14]. However, such hybrid approaches significantly compromise the primary advantage of ML-based surrogate models, their real-time prediction capability, as demonstrated by insufficient speedup factors: for example, approximately 1.9 times acceleration compared to pure computational fluid dynamics (CFD) simulations [14]. This highlights the critical need for improving the long-term prediction accuracy of purely data-driven AR models while preserving their computational efficiency, enabling real-time predictions without relying on expensive numerical solver re-calibrations.
To enhance the robustness of AR-based spatio-temporal predictions over long-term rollouts, the most frequently used approach is noise injection, where noise is added to the input data during the training phase [12, 15, 16, 17]. This method adds random noise during training to help the model become robust against prediction errors that accumulate during the rollout phase. However, noise injection requires careful tuning of the noise scale, which is highly data-dependent, and its stochastic nature can lead to inconsistent and unstable training process.
This study aims to present a novel framework for enhancing long-term AR predictions through the integration of numerical time-integration schemes and adaptive multi-step rollout techniques. Our approach specifically addresses the need for robust long-term prediction methods that maintain the real-time prediction capabilities of data-driven models while performing effectively under practical constraints such as limited datasets and computational resources. Extensive experiments, including comparisons with traditional noise injection methods and evaluations under challenging mesh conditions, have validated the model’s ability to achieve an optimal balance between prediction robustness and computational efficiency. Our key contributions can be summarized as follows:
  1. 1.
    Assessment under challenging but practical conditions: In real-world engineering scenarios, models often encounter challenging constraints such as limited availability of training data, restricted model capacity, and the need to handle varying mesh topologies arising from diverse shape configurations. To address these challenges, we aim to develop a GNN-based AR model that achieves accurate long-term predictions of future states (up to 350 rollouts) using only 50 past snapshots. Our approach emphasizes efficiency by utilizing a lightweight architecture with only 1,177 trainable parameters, specifically focusing on GNNs due to their inherent mesh-scalability and adaptability to diverse topologies.
  2. 2.
    Development of a comprehensive AR framework: We developed a comprehensive AR SciML framework that integrates traditional numerical schemes with advanced multi-step rollout approaches. We systematically evaluated multiple numerical schemes including forward Euler, first- and second-order central differences, and two-step Adams-Bashforth methods. Through extensive testing of these schemes coupled with different multi-step rollout strategies, we achieved significant improvements in prediction accuracy and robustness.
  3. 3.
    Enhancement through adaptive weighting strategies: To further improve the conventional multi-step rollout, we introduced three adaptive weighting strategies. Compared to the standard multi-step rollout with fixed weights, two of these strategies–—with learnable parameters–—demonstrated superior performance. Ultimately, combining the Adams-Bashforth method with an adaptive weighting approach that emphasizes only the first and last future components in the multi-step rollout process enabled accurate prediction of vortex shedding dynamics across all 350 rollouts.
  4. 4.
    Comparison with noise-injection methods: We validated our fusion of the Adams-Bashforth method with adaptive multi-step rollout against the conventional noise injection approach, commonly used to improve long-term rollout performance. Across various time schemes, our adaptive multi-step rollout consistently outperformed noise-injection methods.
  5. 5.
    Robustness evaluation on a more challenging mesh: To assess the robustness of our AR framework with GNN surrogate models, we conducted tests under even more challenging conditions using a truncated mesh that captures only the region behind the cylinder. The combination of the Adams-Bashforth method with adaptive multi-step rollout again significantly outperforms the conventional direct prediction AR approach with vanilla multi-step rollout.
  6. 6.
    General applicability of the proposed approach: Our versatile framework is designed to work beyond specific SciML models, such as GNNs, and specific applications, such as flow-field prediction. It provides a scalable and versatile solution applicable to any AR model across engineering disciplines dealing with spatio-temporal prediction tasks.
The remainder of this paper is organized as follows. Section 2 presents the theoretical framework of our proposed time integration schemes and adaptive multi-step rollout strategies. Section 3 details the overall experimental setup. Sections 4 and 5 demonstrate the effectiveness of time integration schemes and multi-step rollout techniques, respectively. Section 6 introduces our adaptive weighting approaches and analyzes their impact on prediction robustness. Section 7 provides comparative analysis with existing noise injection methods. Section 8 evaluates model performance under more challenging mesh conditions with truncated domains. Finally, Section 9 concludes the paper with discussions on broader implications and future research directions.
2 Methods: time integration schemes and adaptive multi-step rollout
2.1 Auto-regressive rollout with different time integration schemes
2.1.1 Motivation
In the realm of CFD, accurately predicting the temporal evolution of flow fields, particularly over extended time horizons, is crucial for understanding complex flow phenomena. While numerous advanced time-stepping schemes have been rigorously developed and studied in traditional CFD to enhance long-term prediction stability, their application in AI-driven flow prediction remains relatively unexplored. Most AI-based temporal flow prediction research has focused on direct prediction [1, 3, 4, 17] or simple forward Euler methods [2, 15, 16] for temporal integration, often struggling with error accumulation during extended AR rollouts. Consequently, the potential impact of other finite difference schemes on long-term prediction accuracy has largely been overlooked in this context. To address this gap, we propose pioneering the application of various numerical finite difference schemes to enhance long-term AR temporal prediction. Specifically, we integrate and systematically compare discretization techniques commonly used in finite difference methods to understand their behavior, strengths, and limitations within an AR framework, leveraging Graph U-Nets for fluid dynamics predictions over extended time horizons. This investigation represents the first comprehensive study to evaluate the stability and accuracy of diverse time integration schemes in the context of AR temporal prediction for fluid dynamics.
2.1.2 Investigated time integration schemes
In this study, we explore four different finite difference schemes for the time integration of Graph U-Net architecture and compare their performance with the conventional direct prediction approach. The objective is to assess how various finite difference numerical methods impact the accuracy and stability of the temporal evolution of the flow field. Note: The Runge-Kutta method, one of the most widely used time integration schemes in the CFD domain, is not included in this study due to its inherent incompatibility with the data-driven AR process. Further details are provided in A.
0) Conventional approach: direct prediction
In the direct prediction approach, the AR model directly predicts the flow field at the next time step based on input from previous time steps:
$\mathbf{u}(t+\Delta t)=\text{AR}\left(\underbrace{\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)}_{\text{past $N$ snapshots}}\right)$ (1)
where $\mathbf{u}(t)$ represents the flow field at time $t$, $\Delta t$ is the time step size, $N$ is the number of past time steps used as input, and AR refers to the AR model with past $N$ snapshots as inputs. Here, it is important to note that $\Delta t=1$ can be assumed since $\Delta t$ represents the conceptual difference between successive snapshot datasets rather than an actual physical time interval [16]—therefore, $\Delta t=1$ is assumed for the simplicity throughout this study. Returning to the main point, this conventional approach does not explicitly leverage past snapshot datasets, even though they are used as inputs to black-box AR models, ultimately reducing both training and inference efficiency.
1) Forward Euler method
Instead of predicting the next time step directly, forward Euler method [15, 18] makes the AR model predict the rate of change of the flow field:
$\mathbf{f}(t)=\text{AR}\big{(}\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)\big{)}$ (2)
where the rate of change $\mathbf{f}(t)$ is computed by:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}$ (3)
and therefore the next time step can be obtained as:
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\mathbf{f}(t)$ (4)
During training, the label for $\mathbf{f}(t)$ is computed from the given data using the finite difference approximation shown in Eq. 3, and AR model is trained to output the corresponding $\mathbf{f}(t)$ as in Eq. 2. During inference, the AR model predicts $\mathbf{f}(t)$ based on past flow fields, and finally $\mathbf{u}(t+\Delta t)$ can be obtained as in Eq. 4.
2) First-order central difference
With this scheme, the AR model again predicts the $\mathbf{f}(t)$ as in Eq. 2 but the term $\mathbf{f}(t)$ is defined with the form of first-order central difference as [19]:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t-\Delta t)}{2\Delta t}$ (5)
and therefore the next time step can be obtained as:
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t-\Delta t)+2\Delta t\cdot\mathbf{f}(t)$ (6)
This method considers both forward and backward time steps for the calculation of $\mathbf{f}(t)$, potentially offering improved accuracy when flow properties vary smoothly over time. However, when the flow exhibits oscillatory behavior, this approach can lead to erratic predictions. Also, from Eq. 6, note that the next time step $\mathbf{u}(t+\Delta t)$ is highly dependent on $\mathbf{u}(t-\Delta t)$ without the intermediate time step $\mathbf{u}(t)$, which will be discussed later (Fig. 4 in Section 4.1).
3) Second-order central difference
Similar to the first-order central difference, this approach predicts $\mathbf{f}(t)$ but with a different definition using the second-order accurate central difference scheme [20]:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-2\cdot\mathbf{u}(t)+\mathbf{u}(t-\Delta t)}{\Delta t^{2}}$ (7)
and therefore the next time step can be calculated as:
$\mathbf{u}(t+\Delta t)=2\cdot\mathbf{u}(t)-\mathbf{u}(t-\Delta t)+\Delta t^{2}\cdot\mathbf{f}(t)$ (8)
Here, $\mathbf{f}(t)$ in the second-order scheme is defined symmetrically around the time step $\mathbf{u}(t)$ as in Eq. 7 using all three components: $\mathbf{u}(t-\Delta t)$, $\mathbf{u}(t)$, and $\mathbf{u}(t+\Delta t)$. Therefore, it can be expected to reduce the bias in the approximation, making it more reliable and often leading to more stable solutions than the first-order central difference, where there are only two components, $\mathbf{u}(t-\Delta t)$ and $\mathbf{u}(t+\Delta t)$.
4) Adams-Bashforth with forward Euler (Adams-Euler)
Next, we explore the two-step Adams-Bashforth method, which requires derivative information for prediction. The general form of the Adams-Bashforth update is:
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\Delta t\cdot\left(\frac{3}{2}\mathbf{f}(t)-\frac{1}{2}\mathbf{f}(t-\Delta t)\right)$ (9)
where $\mathbf{f}(t)$ is the estimated derivative at time $t$. The method for computing the derivative, $\mathbf{f}(t)$, can be freely chosen by the user. In this study, we focus primarily on the forward Euler method for approximating the derivative $\mathbf{f}(t)$—the Adams-Bashforth with forward Euler method will be referred to as Adams-Euler for the rest of the paper. With this regard, AR model aims to predict the corresponding $\mathbf{f}(t)$:
$\mathbf{f}(t)=\frac{\mathbf{u}(t+\Delta t)-\mathbf{u}(t)}{\Delta t}=\text{AR}\left(\underbrace{\mathbf{u}(t),\mathbf{u}(t-\Delta t),\dots,\mathbf{u}(t-(N-1)\Delta t)}_{\text{past $N$ snapshots from $t$}}\right)$ (10)
Similarly, the previous derivative $\mathbf{f}(t-\Delta t)$ is computed as:
$\mathbf{f}(t-\Delta t)=\frac{\mathbf{u}(t)-\mathbf{u}(t-\Delta t)}{\Delta t}=\text{AR}\left(\underbrace{\mathbf{u}(t-\Delta t),\mathbf{u}(t-2\Delta t),\dots,\mathbf{u}(t-N\Delta t)}_{\text{past $N$ snapshots from ($t-\Delta t$)}}\right)$ (11)
Since both $\mathbf{f}(t)$ and $\mathbf{f}(t-\Delta t)$ can be obtained through Eq. 10 and Eq. 11, the next snapshot can be computed as in Eq. 9. A key feature of the two-step Adams-Bashforth method is its reliance on derivative information from previous time steps, requiring the storage of past outputs ($\mathbf{f}(t-\Delta t)$) to compute future predictions. Due to its property, three successive snapshots, $\mathbf{f}(t-\Delta t)$, $\mathbf{f}(t)$, and $\mathbf{f}(t+\Delta t)$, can be explicitly considered for the prediction of the next snapshot. This approach is particularly well-suited for problems where incorporating derivative approximations from the past is essential for accurately capturing the system’s temporal evolution.
2.2 Proposal of adaptive multi-step rollout
2.2.1 Motivation for multi-step rollout
Conventional AR models trained with the loss function of single-step prediction often struggle with error accumulation during long-term rollouts. In single-step training, the model learns to predict only one step ahead using ground truth data as input, but during inference, it must use its own predictions as inputs for subsequent steps. This creates a mismatch between training and inference conditions, as the model is never exposed to its own prediction errors during training. As a result, when small errors occur in early predictions, they can compound and amplify over rollout processes, leading to significant deviations from the true trajectory in long-term predictions. This points out the necessity of the proposing technique for aiding the AR models to implement well in the long-term rollout tasks, which is covered in details in Section 2.2.2.
2.2.2 Background on conventional multi-step rollout
The multi-step rollout approach (Fig. 1), proposed by Wu et al. [21], addresses the limitation discussed in Section 2.2.1 by incorporating multiple future time steps into the training process. By training the model to predict several steps ahead while iteratively using its own predictions as inputs, this approach enhances robustness against error accumulation. It aligns the training and inference conditions, enabling the model to adapt to and mitigate its own prediction errors.
Specifically, given input snapshots $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, the AR model first predicts $\hat{\mathbf{u}}(t+1)$. The input history is then updated to $\hat{\mathbf{u}}(t+1),\ldots,\mathbf{u}(t-(N-2))$, allowing the model to predict $\hat{\mathbf{u}}(t+2)$. This iterative process continues for $M$ steps, producing predictions from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$ based on the initial input history $\mathbf{u}(t),\ldots,\mathbf{u}(t-(N-1))$, without any backpropagation during inference. The predicted future snapshots, from $\hat{\mathbf{u}}(t+1)$ to $\hat{\mathbf{u}}(t+M)$, are then used to compute the loss: from $\mathcal{L}_{1}$ to $\mathcal{L}_{M}$. In the training process, this method minimizes a multi-step loss function, which aggregates losses across all prediction steps. By explicitly optimizing across multiple steps, the model enhances its resilience to compounding errors during inference. Consequently, the multi-step loss function is defined as a weighted sum of individual step losses, expressed as:
$\mathcal{L}_{multi-step}=\sum_{i=1}^{M}{w_{i}\mathcal{L}_{i}}$ (12)
where $w_{i}$ represents the weight assigned to the loss at each future time step $t+i$. Notably, when $M=1$, this corresponds to the conventional loss without employing the multi-step rollout technique. Backpropagation with respect to $\mathcal{L}_{\text{multi-step}}$ enables the model to optimize for long-term predictions by accounting for all loss components, from $\mathcal{L}_{1}$ to $\mathcal{L}_{M}$.
Figure 1: Overview of multi-step rollout: the model predicts $M$ future steps during training, with the total loss computed as a weighted sum of individual prediction losses at each step. In this study, the adaptive weighting schemes for dynamically determining weights ($w_{i}$) will be newly proposed in Section 2.2.3.
However, this technique presents two main challenges:
Disadvantage 1
While Wu et al. [21] proposed a weighting strategy for each loss term in multi-step rollout, setting $w_{1}=1$ for the first future step and $w_{i}=0.1$ for subsequent steps, these weights typically require manual adjustment depending on the specific problem at hand. Optimal weight selection typically involves trial and error, as the ideal values can vary significantly across different datasets and problem types. Reckless weig ht tuning can destabilize training, leading to issues such as overfitting or underfitting, which compromise model stability and generalization. This instability undermines the effectiveness of multi-step rollout methods, rendering them ineffective for reliable predictions.
Disadvantage 2
The conventional multi-step rollout considers all future $M$ steps simultaneously during training. This can complicate optimization, as the model must account for every snapshot in the long-term future prediction process. As a result, the model might struggle to balance short-term and long-term accuracy, particularly in early stages of training when long-term predictions are highly uncertain. In summary, the simultaneous consideration of all $M$ steps can lead to poor convergence, as errors from long-term predictions dominate, making it harder for the model to achieve stable optimization.
In conclusion, while the conventional multi-step rollout method proposed by Wu et al. [21] is effective for capturing sequential dependencies in time-series data, it presents challenges. Specifically, it requires careful weight tuning and faces the difficulty of optimizing the model simultaneously for all future steps, which can impede convergence and reduce robustness.
2.2.3 Adaptive multi-step rollout
To overcome the limitations of conventional multi-step rollouts, we propose a series of adaptive multi-step rollout strategies designed to enhance training efficiency and improve prediction accuracy over long time horizons, while preserving the scalability of the traditional approach. We introduce three progressively refined ideas to enhance the model’s performance in multi-step predictions. They are named as AW1, AW2, and AW3, where AW refers to the adaptive weighting.
AW1: adaptive weighting without learnable parameter
The first approach seeks to mitigate the need for manually selecting weights $\lambda_{i}$ by automatically assigning weights based on the magnitude of the prediction errors at each time step. Specifically, the weights are computed as the normalized mean squared error (MSE) for each prediction step:
$w_{i}=\frac{\text{MSE}_{i}}{\sum_{j=1}^{M}\text{MSE}_{j}}$ (13)
where $\text{MSE}_{i}$ is the MSE at time step $i$, and $M$ is the total number of rollout steps considered in multi-step rollout process. The total loss is then calculated as:
$\mathcal{L}=\sum_{i=1}^{M}w_{i}\cdot\text{MSE}_{i}$ (14)
This method inherently emphasizes time steps with less accurate predictions, as larger errors yield higher weights. However, because the weights are simply proportional to each MSE value, the approach may lack flexibility in adjusting the importance of different time steps, potentially limiting the model’s capacity to prioritize the most challenging predictions.
Algorithm 1 AW1: adaptive weighting without learnable parameter
1:Model $\mathcal{M}$, input data $\mathbf{u}(t),\dots,\mathbf{u}(t-(N-1))$, ground truth $\mathbf{u}(t+1),\dots,\mathbf{u}(t+M)$
2:Updated model parameters
3:for each training iteration do
4:     Perform $M$-step rollout predictions recursively:
$\hat{\mathbf{u}}(t+1)$ $=\mathcal{M}(\mathbf{u}(t),\dots,\mathbf{u}(t-(N-1)))$
$\hat{\mathbf{u}}(t+2)$ $=\mathcal{M}(\hat{\mathbf{u}}(t+1),\dots,\mathbf{u}(t-(N-2)))$
[…]
$\hat{\mathbf{u}}(t+M)$ $=\mathcal{M}(\hat{\mathbf{u}}(t+M-1),\dots,\mathbf{u}(t))$
5:     Compute MSE for each step:
$\text{MSE}_{i}=\left|\hat{\mathbf{u}}(t+i)-\mathbf{u}(t+i)\right|^{2},\quad i=1,\dots,M$
6:     Compute adaptive weights:
$w_{i}=\frac{\text{MSE}_{i}}{\sum_{j=1}^{M}\text{MSE}_{j}}$
7:     Calculate total loss:
$\mathcal{L}=\sum_{i=1}^{M}w_{i}\cdot\text{MSE}_{i}$
8:     Update model parameters via backpropagation of $\mathcal{L}$
9:end for
10:
11:return Updated model parameters
AW2: adaptive weighting with learnable parameter
To introduce more flexibility in adjusting the importance of different time steps, we extend the first idea by incorporating a learnable parameter $k$ that dynamically adjusts the weighting scheme. The adaptive weights are computed using a power function of the MSE values, modulated by the effective parameter $k_{e}$:
$w_{i}=\frac{\text{MSE}_{i}^{k_{e}}}{\sum_{j=1}^{M}\text{MSE}_{j}^{k_{e}}}$ (15)
where $k_{e}$ is defined as:
$k_{e}=0.5+2.5\cdot\sigma(sk)$ (16)
Here, $\sigma$ represents a sigmoid activation function, used to bound the value of $k_{e}$ within the range of 0.5 to 3. The parameter $k$ is learnable, and $s$ acts as a scaling factor ($s=10$ is adopted in this study). By learning $k$ during training, the model can adjust the weighting more flexibly than AW1, depending on what minimizes the total loss:
$\mathcal{L}=\sum_{i=1}^{M}w_{i}\cdot\text{MSE}_{i}$ (17)
Algorithm 2 AW2: adaptive weighting with learnable parameter
1:Model $\mathcal{M}$, input data $\mathbf{u}(t),\dots,\mathbf{u}(t-N+1)$, ground truth $\mathbf{u}(t+1),\dots,\mathbf{u}(t+M)$
2:Updated model parameters and learnable parameter $k$
3:Initialize learnable parameter $k$
4:for each training iteration do
5:     Perform $M$-step rollout predictions recursively:
$\hat{\mathbf{u}}(t+1)$ $=\mathcal{M}(\mathbf{u}(t),\dots,\mathbf{u}(t-(N-1)))$
$\hat{\mathbf{u}}(t+2)$ $=\mathcal{M}(\hat{\mathbf{u}}(t+1),\dots,\mathbf{u}(t-(N-2)))$
[…]
$\hat{\mathbf{u}}(t+M)$ $=\mathcal{M}(\hat{\mathbf{u}}(t+M-1),\dots,\mathbf{u}(t))$
6:     Compute MSE for each step: $\text{MSE}_{i}=|\hat{\mathbf{u}}(t+i)-\mathbf{u}(t+i)|^{2}$ for $i=1,\dots,M$
7:     Calculate $k_{e}=0.5+2.5\cdot\text{sigmoid}(sk)$
8:     Compute adaptive weights:
$w_{i}=\frac{\text{MSE}_{i}^{k_{e}}}{\sum_{j=1}^{M}\text{MSE}_{j}^{k_{e}}}$
9:     Calculate total loss: $\mathcal{L}=\sum_{i=1}^{M}w_{i}\cdot\text{MSE}_{i}$
10:     Update model parameters and learnable parameter $k$ through backpropagation of $\mathcal{L}$
11:end for
12:
13:return Updated model parameters and $k$
AW3: simplified adaptive weighting focusing on first and last steps
Recognizing that errors often accumulate over time in AR predictions, we propose a simplified version of the second idea (AW2) that only considers the losses on the first and last steps of the rollout. Specifically, the total loss is computed using only the MSE at the first step ($\text{MSE}_{1}$) and the final step ($\text{MSE}_{M}$), ignoring the losses at intermediate steps. In this situation, the adaptive weights are computed as:
$\begin{split}w_{1}=\frac{\text{MSE}_{1}^{k_{e}}}{\text{MSE}_{1}^{k_{e}}+\text{MSE}_{M}^{k_{e}}},\\ w_{M}=\frac{\text{MSE}_{M}^{k_{e}}}{\text{MSE}_{1}^{k_{e}}+\text{MSE}_{M}^{k_{e}}}\end{split}$ (18)
where the two weights are summed to 1. By focusing on the first and last steps, this method emphasizes improving both the immediate next prediction and the long-term prediction, which are critical for model performance. This simplified approach reduces computational complexity and avoids potential overfitting to intermediate steps, while still leveraging the adaptability introduced by the learnable parameter $k$.
Finally, the progression from the first to the third approaches for the adaptive weighting can be summarized as below:
  • AW1 introduces automatic weighting based on MSE, providing simplicity but with limited flexibility due to the absence of additional parameters.
  • AW2 incorporates a learnable parameter that flexibly adjusts the weighting scheme.
  • Aw3 simplifies the AW2 method by focusing on the most critical steps (first and last), reducing complexity while retaining the benefits of adaptivity.
3 Details of the experiments to evaluate spatio-temporal AR prediction performance
In this section, we outline the training process of the Graph U-Net models [22, 17] designed for spatio-temporal flow prediction, specifically applied to the vortex shedding phenomenon behind a two-dimensional (2D) circular cylinder. This specific adoption of a GNN-based model aims to evaluate the performance of our AR framework within a mesh-agnostic SciML paradigm.
3.1 Dataset and preprocessing
[…]
For training, we utilize only 50 consecutive $x$-velocity snapshots from the dataset. Specifically, the model input consists of a sequence of $N=20$ past $x$-velocity snapshots, fewer than the number of snapshots in a vortex shedding period, forming a 20-channel input graph. The task is to predict the flow field of the next immediate snapshot, enabling the model to learn temporal dependencies within the limited training data.
Figure 2: Mesh used for training, containing 1,946 nodes, 11,208 edges, and 3,658 volume cells. The flow is from left to right.
3.2 Model architecture
The Graph U-Net architecture employed in this study is designed with a focus on computational efficiency and simplicity, comprising only 1,177 trainable parameters (Fig. 3). The encoder part of the model consists of four graph convolutional network (GCN) layers with channel dimensions decreasing from 20 to 1 in the sequence 20, 15, 10, 5, and 1. A pooling ratio of 0.6 is applied using the gPool layer [22] after each GCN layer in the encoder, as this ratio was found to be optimal in our previous work [17]. This pooling operation reduces the number of nodes by 60% at each step, effectively capturing multi-scale features while maintaining computational tractability.
Figure 3: The Graph U-Net architecture consists of an encoder and a decoder, with skip connections facilitating the transfer of information from the encoder to the decoder.
The decoder mirrors the encoder with four GCN layers, each having a channel dimension of 1. Skip-connections between corresponding layers of the encoder and decoder are incorporated to facilitate the flow of information and improve reconstruction accuracy. The unpooling operations restore the graph to its original size, ensuring that the output mesh has the same dimensionality as the input.
3.3 Training procedure
The model is trained for 5,000 epochs using the Adam optimizer with an initial learning rate of $10^{-3}$. To ensure robustness and capture variability, each training process is repeated three times under identical settings, with the mean performance metrics reported as the final experimental results. During training, the MSE between predicted and ground-truth snapshots serves as the loss function.
3.4 Evaluation and rollout process
After training, the model is evaluated on its ability to predict 350 future snapshots in an AR rollout manner. In each rollout step, the model uses the most recent 20 snapshots to predict the next snapshot. This process is repeated iteratively to generate long-term predictions over the 350-snapshot horizon.
Considering the small model capacity (1,177 parameters) and the limited size of the training dataset (50 snapshots), the rollout process presents a challenging scenario for the model. The extended prediction horizon of 350 snapshots is substantially longer than the training sequence and encompasses multiple periods of vortex shedding. This harsh evaluation setting is intended to test the model’s ability to generalize temporal dynamics and maintain stability over long-term predictions.
The performance of the model is assessed using the averaged MSE over all 350 predicted snapshots:
$\text{MSE}=\frac{1}{S}\sum_{s=1}^{S}\left(\frac{1}{N}\sum_{n=1}^{N}\left(y_{n,s}-\hat{y}_{n,s}\right)^{2}\right),$ (19)
where $\hat{y}_{n,s}$ and $y_{n,s}$ are the predicted and ground-truth values at node $n$ for snapshot $s$, respectively. $N$ is the total number of nodes, while $S$ is the number of future snapshots evaluated during the multi-step rollout, which is set to 350 in this study.
3.5 Underlying reason for harsh conditions
The model is evaluated under intentionally harsh conditions, including limited model capacity (1,177 parameters), minimal training data (only 50 snapshots), and a long prediction horizon of 350 future snapshots. These constraints simulate practical engineering scenarios where only small datasets are available and computational infrastructure for model training is limited. This challenging setup rigorously tests the model’s robustness and its ability to manage error accumulation over time, particularly for complex temporal patterns like vortex shedding. By evaluating under these conditions, we aim to demonstrate the potential of our approach to develop efficient and accurate AR models for spatio-temporal flow prediction, especially in resource-constrained engineering applications.
4 Application of time schemes into auto-regressive GNNs
In this section, we present the results of applying the four different time integration schemes, along with the conventional direct prediction approach, as described in Section 2.1.2. The objective is to assess their own impact on the long-term prediction accuracy and stability of the Graph U-Net model without yet incorporating the multi-step rollout technique (which will be addressed in later sections).
4.1 Predictive error comparison
We evaluate the performance of each time integration scheme by computing the MSE between the predicted and ground-truth flow fields over the entire spatial domain. The MSE values are averaged over 350 future snapshots, providing a comprehensive assessment of each method’s long-term prediction capability. The results are summarized in Table 1.
Table 1: Comparison of MSE for different time integration schemes over 350 future snapshots. Training time based on NVIDIA 3090 GPU is also shown.
Time Integration Scheme MSE Training time [s]
Direct prediction 0.125 2004
Forward Euler 0.138 1867
First-order central 65.024 1879
Second-order central 0.713 2003
Adams-Euler 0.139 1895
As observed in Table 1, the direct prediction (conventional approach), forward Euler, and Adams-Euler schemes yield relatively low MSE values, seemingly having better prediction accuracy. In contrast, the first-order central difference exhibits significantly higher errors. The training time on the NVIDIA GeForce RTX 3090 24GB GPU is also summarized (all of the experiments conducted in this study is based on the same GPU), indicating that the use of different time integration schemes does not significantly affect the time required to train the model.
Among various time integration approaches, Fig. 4 illustrates the predicted snapshots when using the first-order central difference scheme. Four consecutive snapshots after 100 rollouts are shown, revealing a highly oscillatory behavior: snapshots after 100 (Fig. 4(a)) and 102 (Fig. 4(c)) rollouts appear similar, while those after 101 (Fig. 4(b)) and 103 (Fig. 4(d)) exhibit similar patterns. This alternating pattern persists beyond 100 rollouts (not shown here), reflecting a fundamental limitation of the first-order central difference method. Specifically, the next time step, $\mathbf{u}(t+\Delta t)$, depends heavily on $\mathbf{u}(t-\Delta t)$, bypassing the intermediate state at $\mathbf{u}(t)$ (as described in Eq. 6). These results align with established findings that the central difference schemes are generally unsuitable for time integration due to their inherent instability, resulting in oscillatory and non-physical solutions.
(a) After 100 rollout steps
(b) After 101 rollout steps
(c) After 102 rollout steps
(d) After 103 rollout steps
Figure 4: Predicted snapshots using the first-order central difference scheme after (a) 100, (b) 101, (c) 102, and (d) 103 rollout steps. Snapshots after 100 and 102 rollouts display similar patterns, while those after 101 and 103 show similar but alternate patterns.
To further illustrate the performance of the time integration schemes, we visualize the predicted flow fields after 100 rollout steps for selected outperforming schemes: direct prediction, forward Euler, and Adams-Euler. Fig. 5 presents the ground truth flow field and the corresponding predictions and error distributions for selected schemes. Here, the ground truth flow field shows the specific phase during the vortex shedding pattern behind the cylinder. The predictions from the direct prediction, forward Euler, and Adams-Euler methods show noticeable discrepancies from the ground truth, even after only 100 rollout steps. Especially, in terms of direct prediction, it provides time-averaged results, where the flow fields remain constant as the rollout progresses. The corresponding error plots highlight regions with significant prediction deviations, especially in the wake region where complex flow dynamics are most pronounced.
(a) Ground truth flow field at snapshot $t+100$
(b) Direct prediction
(c) Error of Direct prediction
(d) Forward Euler Prediction
(e) Error of forward Euler Prediction
(f) Adams-Bashforth with forward Euler prediction
(g) Error of Adams-Bashforth with forward Euler Prediction
Figure 5: Visualization of the $x$-velocity fields after 100 rollout steps. Although only 100 rollout steps are performed, all the time integration methods yield unsatisfactory results.
4.2 Need for additional techniques
The results indicate that, despite the relatively low MSE values for some schemes, the Graph U-Net models struggle to maintain accurate long-term predictions over extended rollout horizons. The accumulation of errors becomes apparent even within the first 100 rollout steps, well short of the 350-step prediction horizon that is the main objective of this study. The high errors observed in the first-order and second-order central difference schemes suggest that these methods are less robust for this application, potentially due to their reliance on central difference approximations, which may amplify errors in an AR prediction context due to their oscillatory behavior found in Fig. 4.
These findings underscore the limitations of directly applying traditional finite difference schemes within GNN architectures for time integration, revealing insufficient performance over extended prediction horizons. This highlights the need for enhanced techniques to improve stability and accuracy, such as integrating multi-step rollout strategies to address the observed shortcomings in current methods. In the subsequent sections, we will introduce and evaluate the impact of integrating multi-step rollout techniques with the Graph U-Net models. By combining these approaches with appropriate time integration schemes, we aim to enhance the long-term prediction performance and achieve more reliable forecasts over extended time horizons.
5 Extension of time integration schemes into multi-step rollout scenario
In this section, we extend the previously investigated time integration schemes by incorporating multi-step rollout techniques with different values of $M$, where $M$ represents the number of future snapshots considered during training. Specifically, we examine $M=1$, $2$, $4$, and $8$. When $M=1$, it means that the multi-step rollout is not applied, corresponding to the vanilla AR model. By integrating multi-step rollout, we aim to assess whether training the model to predict multiple future steps can enhance long-term prediction accuracy and stability. Note that the weighting strategy suggested by Wu et al. [21] is adopted herein, where the loss weights are set to $w_{1}=1$ for the first future step and $w_{i}=0.1$ for the subsequent steps in the multi-step rollout (see Eq. 12).
5.1 Performance evaluation with multi-step rollout
We evaluate all the time integration schemes with varying values of $M$ and the MSE results for each combination are presented in Table 2. Here, we can see that increasing the multi-step rollout length $M$ introduces significant computational overhead, as evidenced by the increased training times from 1,914s ($M=1$) to 3,131s ($M=8$) shown in Table 2. However, this increased computational cost does not necessarily lead to improved performance. The results reveal a clear distinction between different time integration schemes: While the Adams-Bashforth with forward Euler (Adams-Euler) scheme demonstrates consistently robust performance across all values of $M$ (MSE improving from 0.139 at $M=1$ to 0.070 at $M=4$), other schemes exhibit significant instability. Notably, the direct prediction scheme breaks down completely with NaN errors at $M=4$ and $M=8$, while the forward Euler method shows instability at $M=2$ (MSE of 581.945) and $M=8$ (NaN). Similarly, both first-order and second-order central difference schemes display extremely high MSE values across increasing $M$.
Table 2: MSE comparison for different time integration schemes with varying multi-step rollout lengths $M$. Erroneous values are shown with a gray background.
Time Integration Scheme Number of considered future snapshots
$M=1$ $M=2$ $M=4$ $M=8$
[…]
Adams-Euler 0.139 0.092 0.070 0.071
Averaged time [s] 1914 2086 2633 3131
These results highlight a crucial finding: naively increasing $M$ imposes excessive training constraints on the model, particularly challenging given its limited capacity of only 1,177 parameters. While longer rollout lengths theoretically allow the model to learn longer-term dependencies, they also require the model to maintain stability across more time steps during training. Most time integration schemes struggle with this requirement, leading to degraded performance or complete breakdown. More importantly, these results demonstrate that the selection of an appropriate time integration scheme plays a more critical role in achieving stable AR predictions than the choice of $M$. This is evident from Table 2, where even with optimal $M$ values, most time integration schemes fail to achieve stable predictions, while the Adams-Euler scheme maintains robust performance across all $M$ values.
Rationale behind the superior performance of Adams-Euler time integration
The Adams-Euler scheme, introduced for the first time in this paper for AR prediction, stands out by maintaining stability even at higher $M$ values. This superior performance can be attributed to its ability to leverage information from multiple previous time steps—since three successive snapshots are explicitly considered for the prediction of the next snapshot as in Eq. 9—in a mathematically principled way, enabling stable predictions even with limited model capacity.
To illustrate the improved performance achieved with the Adams-Euler scheme with $M=4$ multi-step rollout (best case in Table 2), we present visual comparisons of the predicted flow fields after 200 and 300 rollout steps: Fig. 6 shows the ground truth and the model’s predictions at these rollout steps. As seen in Figs. 6(a) and 6(b), the prediction at snapshot $t+200$ shows reasonable agreement with the ground truth, capturing the overall flow patterns and vortex shedding behavior at specific phase. This indicates a significant improvement compared to the results in Section 4, where accurate predictions were not achievable even in the 100 rollout predictions.
(a) Ground truth at snapshot $t+200$
(b) Prediction at snapshot $t+200$ (Adams-Euler): MSE=0.091
(c) Ground truth at snapshot $t+300$
(d) Prediction at snapshot $t+300$ (Adams-Euler): MSE=0.105
Figure 6: Visualization of the $x$-velocity fields at future snapshots using the Adams-Euler scheme with $M=4$ multi-step rollout.
However, at snapshot $t+300$ (Figs. 6(c) and 6(d)), the MSE increases from 0.091 at snapshot $t+200$ to 0.105, resulting in a noticeable divergence between the prediction and the ground truth. The accuracy of the flow structures diminishes, with blurred vortex shedding trail. This indicates that while the incorporation of multi-step rollout and the Adams-Euler scheme significantly enhances long-term prediction performance, there remains room for improvement, particularly for prediction horizons extending beyond $t+200$.
6 Application of proposed adaptive multi-step rollout
In this section, we evaluate three adaptive multi-step rollout approaches elaborated in Section 2.2.3, aiming to enhance the robustness and accuracy of long-term predictions by automatically adjusting the loss function weights during training.
Performance of the three adaptive weighting strategies across different time integration schemes, direct prediction, forward Euler, and Adams-Euler, is assessed. Unlike previous sections, the MSE results from this section are calculated based on $x$-velocity data from seven probe points, as shown in Fig. 7. These results are summarized in Table 3. For the direct prediction case, while the vanilla approach with fixed weights (MSE of 0.011) outperforms its adaptive variants, it still falls short compared to the best results achieved with forward Euler and Adams-Euler schemes. Notably, both our proposed approaches (AW2 and AW3, which are adaptive weightings with a learnable parameter) demonstrate strong performance with the forward Euler scheme (MSE of 0.007 and 0.010 respectively), which is a widely adopted approach in GNN-based AR prediction. However, the most remarkable performance is achieved by combining the Adams-Euler scheme with AW3, which emphasizes only the first and last loss components during training. This combination yields a significantly lower MSE (0.002) compared to all other methods, establishing it as the optimal approach for long-term prediction in our study. A comparison of the required training time reveals negligible differences across the various weighting approaches. This indicates that the proposed adaptive weighting methods enhance rollout performance without incurring additional computational costs.
Figure 7: Locations of seven probe points examined to gain a more intuitive understanding of vortex shedding prediction in terms of $x$-velocity.
Table 3: MSE comparison for different adaptive weighting approaches across time integration schemes: multi-step rollout with $M=4$ is applied. MSE values here are calculated with respect to seven probe points in Fig. 7.
Time Scheme Vanilla AW1 AW2 AW3
(fixed weights) (without learnable $k$) (with learnable $k$) (first and last only)
Direct prediction 0.011 0.025 0.027 0.027
Forward Euler 0.019 0.023 0.007 0.010
Adams-Euler 0.018 0.017 0.010 0.002
Averaged time [s] 2420 2384 2408 2354
To understand how the adaptive weights are adjusted during training in AW2, we visualize the weights assigned to each loss term over epochs for the Adams-Euler scheme as Fig. 8, which reveals two key observations:
  • Stabilization phase (before 1000 epochs) During the initial training phase, the Graph U-Net stabilizes, with minimal differences observed between the weights of the four loss terms. This indicates that the model is primarily learning basic temporal patterns, while the adaptive weighting mechanism has not yet significantly shifted its focus toward specific time steps.
  • Adaptation phase (after 1000 epochs) Beyond epoch 1000, the weight assigned to the 4th loss term increases relative to the others. The weights for the first three loss terms do not exhibit clear trends or significant differences among themselves.
Figure 8: Evolution of adaptive weights for each loss term over epochs in AW2 with the Adams-Euler scheme. Adaptive weights ($w_{i}$) and effective parameter ($k_{eff}$) are shown in each y-axis.
This trend supports the rationale behind AW3, which focuses on the first and last loss components. The increasing weight of the last loss term indicates that it is more challenging to minimize and is critical to long-term rollout performance—remember that all AW approaches are designed to give more weight to the higher loss term. The lack of distinction among the first three loss terms suggests that emphasizing the first loss term can sufficiently represent the short-term prediction error, allowing the model to allocate more resources to improving long-term accuracy. The corresponding results of the Fig. 8 with respect to the AW3 approach can be found in Fig. 13 in B.
To intuitively and quantitatively evaluate the impact of the improvements achieved by AW3 on capturing vortex shedding behavior, we analyze the temporal evolution of the $x$-velocity at seven probe points positioned downstream of the cylinder (Fig. 7). Based on these probe points, we compare the predictions of the two best-performing models trained so far (Table 3). Model A: the one with forward Euler scheme and AW2 (adaptive weighting with learnable parameter), resulting in an MSE of 0.007. Model B: the other with Adams-Euler scheme with AW3 (first and last loss terms only), resulting in an MSE of 0.002.
Fig. 9 shows the time-series of the $x$-velocity at the probe points over 350 future snapshots for both models, along with the ground truth. For Model A (Fig. 9(a)), the predictions generally capture the vortex shedding patterns, with periodic oscillations similar to the ground truth. However, discrepancies are noticeable, especially at 4th probe point with red colors, where the model fails to accurately reproduce the amplitude and phase of the oscillations. However, in the results of Model B (Fig. 9(b)), the predictions closely match the ground truth across all probe points, including 4th probe point. The model successfully maintains the correct shedding period and amplitude over the entire prediction horizon, indicating a significant improvement in long-term accuracy. These results demonstrate that focusing on the first and last loss terms during training (AW3) enhances the model’s ability to capture both short-term and long-term temporal dynamics more effectively. This approach yields predictions that are both more accurate and stable compared to AW2, where all loss terms are treated simultaneously.
(a) Model A: Forward Euler with AW2
(b) Model B: Adams-Euler with AW3
Figure 9: Time series of $x$-velocity at the seven probe points over 350 future snapshots: comparison between (a) Model A and (b) Model B. From bottom to top, the plot represents the $x$-velocity at probe points 1 through 7 (Fig. 7). Solid lines represent ground truth; circles represent model predictions.
From Fig. 9(b), we have demonstrated successful long-term prediction over 350 rollout steps using an AR Graph U-Net model, despite operating under notably challenging conditions: a severely constrained model capacity (1,177 parameters), minimal training data (50 snapshots), and an extensive prediction horizon (350 rollouts). This achievement can be attributed to two key innovations: (1) the introduction of the Adams-Euler time integration scheme for AR prediction, which provides enhanced numerical stability, and (2) the development of an adaptive weighting strategy that strategically focuses on only the first and last future snapshots during multi-step rollout training.
To further validate the effectiveness of these innovations, the following sections will present two critical comparisons: first, against the conventional noise injection approach widely used for improving long-term AR predictions (Section 7), and second, by testing our model under even more challenging conditions using a truncated mesh domain (Section 8). These comparisons will serve to conclusively demonstrate the robustness and superiority of our proposed methodology.
7 Comparison with conventional noise-injection approach for long-term rollout
In this section, we compare the long-term rollout performance of our models against the most widely adopted approach in the field of auto-regression: noise injection [12, 15, 16, 17]. This method deliberately adds random noise to the input data during the training phase to enhance the model’s robustness against error accumulation in AR predictions. The underlying principle is that by exposing the model to perturbed input data during training, it becomes more resilient to prediction errors that naturally accumulate during the rollout phase.
We evaluate three model configurations, all trained with Gaussian noise $\mathcal{N}(0,0.16^{2})$ injected into the input data. This noise level, identified as optimal in our previous work [17] on the same dataset, is consistently applied during training. Notably, evaluations are conducted without multi-step rollouts. The first model employs the traditional direct prediction approach, the second uses the forward Euler time integration scheme, and the third adopts the Adams-Euler time integration scheme. During inference, no additional noise is introduced, allowing the models to leverage their learned robustness to maintain performance over extended rollouts.
To assess the effectiveness of noise injection, we again analyze the temporal behavior of the $x$-velocity at seven probe points located downstream of the cylinder, as in previous evaluations (refer to Fig. 9). Figs. 10(a) and 10(b) present the results for the forward Euler and Adams-Euler scheme with noise injection, respectively. Comparing Fig. 10(a) with Fig. 9(a), which adopts forward Euler with multi-step rollout using AW2, we observe that the model with noise injection exhibits inferior accuracy at all probe points. The oscillations captured by the model deviate more significantly from the ground truth in terms of both amplitude and phase, indicating that noise injection alone is insufficient to maintain long-term prediction accuracy. Similarly, comparing Fig. 10(b) with Fig. 9(b), which adopts Adams-Euler and multi-step rollout using AW3, the model with noise injection shows less accurate reproduction of the vortex shedding patterns, especially for the lower four probe points. Although some oscillatory behavior is present, the predictions do not align closely with the ground truth, especially over extended rollout steps after 150 rollouts. Finally, it is worth noting that when comparing Figs. 10(a) and 10(b),the rollout performance of the Adams-Euler time integration scheme significantly surpasses that of the forward Euler approach, even in the case of noise injection. This observation further underscores the robust performance and superior accuracy of the Adams-Bashforth time integration scheme over the simple forward Euler method across various scenarios.
(a) Forward Euler with noise injection
(b) Adams-Bashforth with forward Euler and noise injection
Figure 10: Time series of $x$-velocity at probe points over 350 future snapshots using models with noise injection. Solid lines represent ground truth; circles represent model predictions.
To quantify the performance differences, we summarize the MSE with respect to seven probe points for each model configuration in Table 4. The results indicate that, for all time integration schemes, models trained with adaptive multi-step rollout consistently outperform those using noise injection. Specifically, the MSE for the Adams-Euler scheme improves from 0.012 with noise injection to 0.002 with adaptive multi-step rollout—a substantial enhancement in prediction accuracy. Although noise injection improves training efficiency by eliminating the need for additional processing of multiple output snapshots, its cost benefits are surpassed by the superior accuracy achieved with our proposed approach.
Table 4: MSE comparison between adaptive multi-step rollout and noise injection approaches for different time integration schemes. Each time scheme utilizes a different adaptive weighting approach for multi-step rollout, selected as the best configuration for each case. MSE values here are calculated with respect to seven probe points in Fig. 7.
Time Scheme Multi-Step Rollout Noise Injection
Direct Prediction 0.0111 0.019
Forward Euler 0.0072 0.017
Adams-Euler 0.0023 0.012
Averaged time [s] 2354 1806
  • 1
    Without adaptive weighting
  • 2
    With AW2
  • 3
    With AW3
The performance gap between adaptive multi-step rollout and noise injection can be attributed to fundamental differences in how these methods address error accumulation. Noise injection attempts to build resilience by introducing input perturbations, making the model more robust to small deviations. However, this approach does not explicitly address the temporal dependencies and error propagation mechanisms inherent in AR predictions. In contrast, our adaptive multi-step rollout strategy enables the model to learn from multiple future steps simultaneously, using strategically adjusted loss weights to focus on both immediate and distant predictions. The effectiveness of this approach is clearly demonstrated in the time series visualizations, where models trained with adaptive multi-step rollout more accurately capture the vortex shedding patterns compared to those trained with noise injection, as shown in Figs. 9 and 10. These results establish our proposed methodology, combining Adams-Euler time integration with adaptive multi-step rollout, as a more effective and reliable alternative to the traditional noise injection technique for long-term prediction tasks.
Combination of multi-step rollout and noise injection techniques
Experiments combining multi-step rollout and noise injection were also conducted; however, their performance fell significantly short compared to methods utilizing either multi-step rollout or noise injection individually. The results suggest that the simultaneous use of these long-term rollout techniques can degrade model performance. This degradation is likely due to the excessive noise introduced into the input datasets during training, stemming both from the artificial noise added through the noise injection approach and the predictive noise generated from the multi-step rollout outputs of the AR models.
8 Robustness evaluation under more challenging conditions: training with truncated mesh regions
Having achieved successful long-term predictions under stringent conditions—limited datasets, small model capacity, and extensive rollout steps—using the novel integration of the Adams-Bashforth with forward Euler scheme and multi-step rollout with adaptive weighting (Fig. 9(b)), we further challenge the model’s robustness by evaluating its performance on a truncated mesh. This truncated mesh contains only a portion of the vortex shedding region, specifically defined by the spatial domain $0.750.3<x<$ 0.75 and $0.3280.128<y<$ 0.328. By focusing on this core region where complex flow dynamics occur, we aim to assess the model’s ability to generalize and maintain accuracy under even challenging conditions.
8.1 Experimental setup for truncated mesh scenario
The truncated mesh is a subset of the original computational domain, capturing the essential features of the vortex shedding phenomenon while reducing the spatial extent of the data (region of interest within the flow field is visualized in Fig. 11). We evaluate the performance of the four models that previously demonstrated satisfactory results, as reported in Table 3: these models include the direct prediction with vanilla multi-step rollout, forward Euler with adaptive weighting approach 2 (AW2), forward Euler with adaptive weighting approach 3 (AW3), and Adams-Euler combined with adaptive weighting approach 3 (AW3).
8.2 Results and discussion
Figs. 11(b)$\sim$11(e) present the predicted flow fields after 300 rollout steps from the four model settings mentioned above, alongside the ground truth in Fig. 11(a). We observe that the direct prediction (Fig. 11(b)) with vanilla multi-step rollout fails to capture the vortex shedding patterns, showing significant deviations from the ground truth. The forward Euler method with AW2 (Fig. 11(c)) shows some improvement but still struggles to accurately represent the flow structures, especially in the wake region behind the cylinder. The forward Euler method with AW3 (Fig. 11(d)) exhibits better performance than the previous two models, capturing some features of the vortex shedding but with noticeable discrepancies. In contrast, the Adams-Euler scheme combined with AW3 (Fig. 11(e)) closely matches the ground truth, successfully reproducing the vortex shedding patterns and flow dynamics within the truncated region. These results highlight the superior performance of the Adams-Bashforth with forward Euler scheme combined with adaptive weighting approach 3, even under the harsher condition of a truncated mesh.
(a) Ground truth flow field at snapshot $t+300$
(b) Prediction by direct prediction with vanilla multi-step rollout
(c) Prediction by Forward Euler with AW2
(d) Prediction by Forward Euler with AW3
(e) Prediction by Adams-Euler and AW3
Figure 11: Comparison of predicted $x$-velocity fields in the truncated-mesh region after 300 rollout steps. Note that only the part of the mesh shown in the figures was used for training Graph U-Nets.
To further assess the models’ ability to capture temporal dynamics within the truncated mesh, we analyze the temporal evolution of the $x$-velocity at seven probe points strategically located in the wake region, as illustrated in Fig. 12. The results demonstrate clear performance differences across the four model configurations. The direct prediction method (Fig. 12(a), MSE: 0.019) completely fails to reproduce the periodic oscillations characteristic of vortex shedding, indicating its fundamental inability to capture flow dynamics in the truncated domain. The forward Euler method with AW2 (Fig. 12(b), MSE: 0.011) shows improved performance, successfully capturing the overall oscillatory behavior, though it gradually deviates from the ground truth in both amplitude and phase as the rollout progresses. Similarly, the forward Euler method with AW3 (Fig. 12(c), MSE: 0.013) achieves comparable accuracy to AW2, despite of the slightly lower performance. Most notably, the Adams-Euler scheme combined with AW3 (Fig. 12(d), MSE: 0.008) demonstrates superior performance, achieving the best agreement with ground truth across nearly all probe points. This configuration accurately captures the vortex shedding frequency, amplitude, and phase throughout the entire prediction horizon, with some deviations observed at probe point 4 (colored as red). These results further reinforce the robustness of the Adams-Euler scheme with adaptive multi-step rollout, even under the challenging conditions of the truncated mesh.
(a) Direct prediction with vanilla multi-step rollout (MSE=0.019)
(b) Forward Euler with AW2 (MSE=0.011)
(c) Forward Euler with AW3 (MSE=0.013)
(d) Adams-Euler and AW3 (MSE=0.008)
Figure 12: Time series of $x$-velocity at probe points over 350 future snapshots for different models trained on truncated meshes. Solid lines represent ground truth; circles represent model predictions.
8.3 Final remarks on the proposed framework: Adams-Bashforth-based adaptive multi-step rollout
The evaluation under the more challenging condition of a truncated mesh provides compelling evidence for the exceptional robustness of our proposed methodology—the synergistic combination of the Adams-Bashforth with forward Euler scheme and adaptive weighting approach 3. Despite the increased complexity of predicting flow dynamics with reduced spatial information in a confined region, this integrated approach consistently demonstrates superior performance over alternative methods. While the direct prediction and forward Euler methods, even with various weighting strategies, struggle to maintain accuracy in this challenging scenario, our combined approach continues to deliver high prediction accuracy and stability. This noticeable performance contrast highlights a crucial insight: the selection of appropriate time integration schemes, when coupled with effective training strategies, plays a fundamental role in developing robust AR models.
The successful performance of our methodologies under these more challenging conditions—beyond those in our previous experiments—underscores their broad applicability and reliability. This robustness can be particularly powerful in practical engineering applications where data availability is often limited or analysis must focus on specific regions of interest within complex flow fields due to the memory limitations. By delivering accurate predictions even with truncated spatial information, our approach establishes itself as a robust and versatile tool for a wide range of real-world flow prediction scenarios. It offers a compelling solution in situations where traditional methods, such as forward Euler time integration, vanilla multi-step rollout, and noise injection, may fall short or prove insufficient.
9 Conclusions and future work
This study presents a comprehensive framework for enhancing long-term auto-regressive (AR) predictions in SciML models through the novel proposal of numerical time-integration schemes and adaptive multi-step rollout techniques. Our approach addresses the fundamental challenge of error accumulation in AR models while maintaining computational efficiency and broad applicability across various scientific domains. By combining the Adams-Bashforth forward Euler scheme with adaptive multi-step rollout strategies, we have demonstrated significant improvements in both prediction accuracy and stability over extended time horizons.
The quantitative findings obtained through the realization of each key contribution introduced in Section 1 can be summarized as follows:
  1. 1.
    Development of a comprehensive AR framework: While direct application of various time integration schemes showed minimal improvement over conventional direct prediction in long-term rollout performance, their integration with multi-step rollout revealed significant differences. The Adams-Bashforth combined with the forward Euler scheme achieved a 7% improvement in prediction accuracy compared to the commonly used forward Euler schemes, whereas the traditional direct prediction approach failed to converge. While other time integration schemes exhibited instability as the number of future time steps increases in multi-step rollout training, Adams-Bashforth with forward Euler demonstrated consistent and robust performance throughout varying prediction horizons.
  2. 2.
    Enhancement of conventional multi-step rollout through adaptive weighting: The proposed adaptive weighting strategy focusing on first and last components delivered 89% improvement over conventional fixed weights and outperformed other adaptive approaches by up to 80%, while maintaining similar computational costs. This demonstrates the effectiveness of selectively emphasizing critical time steps during training rather than considering all future steps simultaneously during multi-step rollout.
  3. 3.
    Success under challenging conditions for long-term rollout: Simulating practical scenarios where computational resources and data availability are limited, we used only 50 training snapshots and 1,177 model parameters. Under resource-constrained conditions, accurate predictions over 350 time steps were achieved by integrating the Adams-Bashforth and forward Euler schemes with adaptive weighting that prioritized the first and last loss components. This level of performance was unattainable with conventional AR techniques.
  4. 4.
    Validation of the proposed approach over traditional noise injection approach: Our approach achieved an 83% improvement in prediction accuracy compared to the traditional noise injection method, a widely used technique for enhancing the rollout performance of AR models. Besides, we observed a novel finding: combining multi-step rollout with noise injection led to degraded performance. This decline is attributed to the compounding effect of excessive noise arising from both artificial injection and errors accumulated during AR-based multi-step predictions.
  5. 5.
    Robustness in more challenging scenario with truncated mesh: In tests with a truncated mesh domain, our method achieved 58% and 27% improvement over direct prediction and forward Euler approaches respectively. This successful performance with reduced spatial information demonstrates broad applicability for practical engineering scenarios where computational memory constraints often necessitate focusing on specific narrow regions of interest within complex flow fields.
  6. 6.
    General applicability of the proposed approach: Through extensive validation, we have shown for the first time that integrating Adams-Bashforth time-stepping with adaptive multi-step rollout significantly enhances the performance of AR predictions. This versatile framework is compatible with any SciML architecture designed for AR prediction, setting a new standard for long-term temporal forecasting across diverse engineering domains.
While our study marks significant advances in AR prediction capabilities, several opportunities for future research remain. First, integrating more advanced higher-order time integration schemes could further improve prediction stability and accuracy. Second, as multi-step rollout involves computing gradients across multiple future steps, developing more efficient and tactical approaches to reduce its computational overhead would be a valuable direction. Finally, extending the application of our framework to other scientific domains and exploring its performance with diverse neural network architectures and temporal dynamics would provide deeper insights into its broader utility and potential limitations, paving the way for more versatile and scalable solutions.
Acknowledgments
This work was supported by the National Research Foundation of Korea (2018R1A5A7025409), and the Ministry of Science and ICT of Korea grant (No. 2022-0-00969, No. 2022-0-00986, and No. RS-2024-00355857), the Ministry of Trade, Industry & Energy grant (RS-2024-00410810), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-300). R.V. was supported by ERC Grant No. 2021-CoG-101043998, DEEPCONTROL. The views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of European Union or European Research Council.
Appendix A On the applicability of Runge-Kutta methods for AR prediction
While Runge-Kutta (RK) methods, particularly the fourth-order scheme, are widely used in numerical integration due to their superior accuracy, their application to AR prediction presents fundamental challenges that make them impractical for our framework. This section explains why RK methods were not explored in our study despite their popularity in traditional numerical integration.
A.1 Theoretical background of RK
The classic fourth-order Runge-Kutta (RK4) method for solving $dy/dt=\mathbf{f}(t,\mathbf{u})$ requires four evaluations per time step:
$k_{1}$ $=\mathbf{f}(t,\mathbf{u}(t))$ (20)
$k_{2}$ $=\mathbf{f}(t+\Delta t/2,\mathbf{u}(t)+(\Delta t/2)k_{1})$ (21)
$k_{3}$ $=\mathbf{f}(t+\Delta t/2,\mathbf{u}(t)+(\Delta t/2)k_{2})$ (22)
$k_{4}$ $=\mathbf{f}(t+\Delta t,\mathbf{u}(t)+\Delta tk_{3})$ (23)
The solution is then advanced using:
$\mathbf{u}(t+\Delta t)=\mathbf{u}(t)+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4})$ (24)
A.2 Fundamental challenges of RK in AR context
A.2.1 Ground truth availability
The primary obstacle in implementing RK methods for AR prediction is the absence of ground truth data for intermediate stages. While $k_{1}$ can be computed directly from available data as $(\mathbf{u}(t+\Delta t)-\mathbf{u}(t))/\Delta t$, the subsequent stages $k_{2}$ and $k_{3}$ require evaluations at intermediate time points $(t+\Delta t/2)$ for which no data exists in typical spatio-temporal datasets. This presents an critical challenge for supervised training of AR models.
A.2.2 Error propagation
AR prediction inherently suffers from error accumulation, where small prediction errors compound over time as the model uses its own predictions as inputs for subsequent steps. The integration of RK methods would significantly exacerbate this fundamental limitation through a multi-stage compound effect:
  • Initial errors in $k_{1}$ prediction affect the computation of intermediate state for $k_{2}$
  • These amplified errors then propagate to $k_{3}$ computation through another intermediate state
  • Errors further accumulate in $k_{4}$ calculation using the previous error-affected states
  • All these cascaded errors combine in the final state prediction, which then feeds into the next AR step
This creates a “double cascade” of errors: first through the RK stages within each time step, and then through the AR prediction chain across time steps. While traditional RK methods are designed to reduce numerical integration errors in classical differential equations, their multi-stage nature paradoxically amplifies the inherent error accumulation problem in AR prediction. This dual error amplification makes the training process inherently unstable and potentially divergent, rendering RK methods impractical for AR applications despite their popularity in CFD domain.
A.2.3 Computational considerations
The RK4 method would require four separate model evaluations per time step, significantly increasing computational cost compared to our adopted Adams-Bashforth scheme. This increased cost is particularly problematic in the context of long-term predictions where hundreds of time steps may be required.
A.3 Contrast with Adams-Bashforth scheme
The Adams-Bashforth scheme, as implemented in our work, avoids these challenges by:
  • Requiring only directly computable derivatives
  • Utilizing historical information rather than intermediate states
  • Needing only one model evaluation per time step
While RK methods offer superior accuracy in traditional numerical integration contexts, their fundamental requirements make them unsuitable for AR prediction tasks. The lack of ground truth for intermediate stages and the compound nature of prediction errors present critical challenges for practical implementation. These limitations explain the reason for our choice of the Adams-Bashforth scheme, which provides a more practical and robust approach to long-term AR prediction.
Appendix B Evolution of the weights in AW3-based multi-step rollout
In Section 6, the weights for each loss term in the AW2-based multi-step rollout are visualized in Fig. 8. This section extends that analysis by presenting results from the AW3-based approach, using the same model configuration as in Fig. 8 but applying AW3 instead of AW2. Fig. 13 illustrates these results, revealing that after 1,000 epochs, the model begins to stabilize the weights of the first and last future steps, with the last time-step weight slightly exceeding that of the first. This indicates that AW3, which focuses only on the first and last time-step losses, still performs comparably well without the intermediate loss terms required by AW2—indeed, AW3 shows the MSE of 0.002 where AW2 shows 0.010 as in Table 3.
Figure 13: Evolution of adaptive weights for each loss term over epochs in AW3 with the Adams-Euler scheme.